OpenStudy (emmaleelooney):

If someone would help me answer this question, I'll be glad to medal or fan; both. Identify the maximum or minimum value and the domain and range of the graph of the function y = 2(x + 2)2 - 3 A. minimum value: 3 domain: all real numbers 3 range: all real numbers B. maximum value: -3 domain: all real numbers ≤ 3 range: all real numbers C. maximum value: 3 domain: all real numbers range: all real numbers ≤ 3 D. minimum value: -3 domain: all real numbers range: all real numbers ≥ -3

OpenStudy (jango_in_dtown):

\[y=2 (x+2)^2 -3\]

OpenStudy (jango_in_dtown):

??

OpenStudy (jango_in_dtown):

see (x+2)^2>=0

OpenStudy (jango_in_dtown):

so 2(x+2)^2>=0

OpenStudy (jango_in_dtown):

implies 2(x+2)^2-3>=-3

OpenStudy (jango_in_dtown):

so minimum value is -3

OpenStudy (jango_in_dtown):

I guess you can say the answer now.

OpenStudy (jango_in_dtown):

obviously x can take any real value and so the domain is the set of all real numbers,..

OpenStudy (jango_in_dtown):

@emmaleelooney